# THERADIA WoZ: An Ecological Corpus for Appraisal-based Affect Research in Healthcare

### v1

@@ IV Data segmentation and annotation
We collected about 80 hours of data from the 52 healthy older partipants, and about 6 hours of data from the 9 MCI participants. We describe in this section the methods used for the segmentation, transcription, and annotation of those data.
@@ IV-A Segmentation, transcription and selection of sequences
The segmentation of the recordings was based on variations in breath groups, ensuring they form a single proposition if they conveyed the same semantic information, or separate propositions if they related to different types of semantic information. Non-verbal expressions occurring within or proximate to a proposition were marked by a diacritic, and those occurring at distance were specifically segmented. Temporal markers for proposition start and stop were both positioned at a certain distance to allow a sufficient observation space for the affect annotation.
The transcription relied mostly on the ESLO project principles [35], respecting spelling and spoken structures; details regarding our convention are reported in section II.A of the supplementary material. Recordings were transcribed with the ELAN software [36] by five Master students in Language Engineering at the Université Grenoble Alpes with a two-steps process: an initial stage of segmentation and transcription, followed by a subsequent verification stage conducted by another student. The segmented sequences were selected to retain only those containing expressions of affect, and a fraction was randomly removed to reduce the overall duration. Obtained sequences were then all reviewed and some were further removed to reduce redundancy. Statistics of the sequences are given in table I.
@@ IV-B Selection of affective labels and dimensions
The context of CCT aligns closely with the literature on achievement affects [37, 38, 39, 40, 41], which considers the following types of affect: (i) the activity affects – generated by the task achievement itself; (ii) the outcome affects – generated by the success or failure results, whether retrospective or prospective; (iii) the epistemic affects – generated by a form of meta-cognition on the perception of the achievement; and (iv) the social affects – generated by another individual.
These affects are particularly relevant in the context of AI-assisted healthcare, as they comprehensively cover those that can be experienced during CCT, i.e., those generated by the CCT itself (activity affects), those generated by past or future CCT failure or success (outcome affects), those generated by the perception of the cognitive abilities involved in CCT (epistemic affects), and those generated by AI-patients interactions (social affects). Our categorical representations of affect relied on the following 23 labels: angry, annoyed, anxious, ashamed, confident, contemptuous, curious, desperate, disappointed, embarrassed, excited, frustrated, interested, guilty, happy, hopeful, impatient, proud, relaxed, satisfied, sad, surprised, and upset.
@@ IV-C Annotation tool and guidelines
We took benefits from previous collaborations with the company Viadialog [^4] https://viadialog.com/ engaged in the development of a web-based annotation tool originally tailored for audio data: ANNOT. This tool has been adapted to the needs of our project and is available online [^5] The URL will be provided for camera-ready. The platform allowed annotators – 6 French students (3 females) from different backgrounds (engineering, business, and biology) and places (Lyon and Toulouse) – to view and annotate the selected audiovisual sequences at home, on their own computer.
The annotation consisted of three steps: a time-continuous annotation, cf. Figure 3, followed by a summary value, for each of the five dimensions, and a summary value of intensity for one or multiple labels. Thus, each sequence was viewed five times, once for each dimension. At the end of the time-continuous annotation, a graph showing the dynamics of the dimension over time was presented to the annotators, who could then validate or perform the annotation again. After the validation, annotators were asked to provide a summary value for the whole sequence. Finally, once all dimensions were annotated, annotators were asked to indicate if one or more labels could describe the affect they had recognised in the sequence and provide an intensity value for them. Annotators could annotate as many labels as they wished, or none at all if the sequence did not contain any particular affect. Further details on the annotations steps are reported in section II.B of the supplementary material.
@@
A document was provided for each annotator as guidelines. It was explained that the annotation task had to be performed in full screen mode in a silent place. The document advised against annotating too many sequences consecutively and suggested a ten-minute break after an hour of work. We also stressed that there were no right or wrong answers in the annotation, as the information to be judged was subjective in nature. Thus, the annotations were to be made by trying to put oneself in the participant’s shoes to assess his or her feeling about the expression produced in the selected sequences. These expressions could include the chosen words, the way they were communicated orally, and the expressive dynamics of the face and body. Finally, the document provided definitions about dimensions and labels. Each dimension was defined as follows:
@@ V Corpus Analysis
This section provides a comprehensive statistical analysis of the annotation data. We first assessed the annotation frequency of the affective labels as well as the inter-annotator agreement on their frequency and intensity. We then evaluated the cognitive delay involved in the annotation of each dimension, as well as the inter-annotator agreement. Relationships between dimensions and affective labels were also estimated, and a label core set was finally selected, based on compliance with consistency criteria relating to each of the analyses mentioned above.
Figure 4: Frequency of the annotated labels according to the agreement of at least one, two or three annotators among six; the vertical axis is on a logarithmic scale to ease the reading.
@@ V-A Categorical annotations
The frequency at which the affective labels were selected by at least one, two, or three annotators is reported in Figure 4. Based on the sample of annotation frequency per label for at least one annotator, Cohen’s $d$ was calculated to identify the theoretical value at which the effect size of its deviation from the sample mean can be considered at least "small" (i.e., Cohen’s $d$ > 0.2  [43]). This theoretical value was equal to 747 sequences. The annotation frequency per label for at least one annotator is reported in Table II, and shown in bold when above this theoretical value.
@@
where $k$ is the total number of classes, which corresponds here to the presence or absence of a label. $N^{i}$ is the total number of annotated sequences for a given affect label, and $N_{c}^{i}$ is the number of commonly identified classes, whether present or absent. Results are reported in Table II. According to a two-tailored $z$-test, the inter-annotator agreement would be defined as statistically above chance if it would exceed 53%, which is the case for all the labels excepted embarrassed; those labels are marked in bold style in Table II.
TABLE II: Statistics on the annotated affective labels: frequency, inter-annotator agreement on the presence, and intensity, and correlation to appraisal’s dimensions. Most relevant labels in the context of AI-assisted CCT are marked in bold.
Label Frequency Agreement on Correlation to
@@
We further evaluated the inter-rater agreement on the intensity of the affective labels using the PCC. The inter-rater agreement was defined as sufficiently high if its Cohen’s $d$, i.e., its correlation coefficient value divided by the sample standard deviation, would be superior to .2 [43], which corresponds to a PCC greater than .065. This concerned most of the affective labels, which are reported in bold style in Table II.
TABLE III: Inter-annotator agreement on each sequence, or CCT session, computed with the Pearson’s correlation coefficient for both time-continuous and summary values of appraisal dimensions; mean (standard-deviation).
@@ V-B Dimensional annotations
The cognitive delay appearing in the continuous annotation of each dimension was evaluated, and results showed that the delay associated with the annotation of goal conduciveness (3.3 seconds) was significantly longer than the delay observed on the other dimensions (less than 2 seconds), suggesting that annotators had more difficulty assessing it. Further details on this evaluation are provided in section III.A of the supplementary material.
@@
We assessed the inter-annotator agreement on the dimensions for both time-continuous and summary annotations, using the Pearson Correlation Coefficient (PCC), which was computed either on each sequence, or each CCT session, cf. Table III. An high agreement was observed on the dimensions of coping, goal conduciveness, and intrinsic pleasantness, with a more pronounced consistency observed in summary annotations compared to time-continuous annotations, which is expected since time-continuous data imply a greater variation in the possible outcomes. A lower agreement was however found for arousal and novelty, which might be explained by several reasons as detailed below.
@@
The novelty dimension was assumed to be more difficult to annotate, as it probably concerned very specific cases of our study, e.g., unplanned feedback from the virtual assistant on the participant’s performance, occurring in fewer sequences than the other dimensions. This assumption was corroborated by an analysis evidencing that the frequency of high-value annotations was lower for novelty compared to other dimensions; for a high-value threshold set at .85, the proportion of high-value annotations for novelty accounted for only .13% of the time-continuous annotation data, whereas other dimensions were at least twice more frequent: .30% for arousal, .24% for intrinsic pleasantness, .62% for goal conduciveness, and .36% for coping. A similar trend was observed in summary annotations, with the occurrence of high-value annotations for novelty accounting to .26%, whereas all other dimensions were twice as prevalent; arousal: .92%, intrinsic pleasantness: .52%, goal conduciveness: 1.35%, and coping: 2.16%.
Figure 5: Heatmap of the PCC between each dimension (summary value) presented in the columns – AR: arousal, CP: coping, GC: goal conduciveness, IP: intrinsic pleasantness, and NV: novelty – and each affective label on the rows, which are sorted in alphabetical order from positive (+) to negative (-). Significant correlations (p-value $\leq{0.05}$) are coloured either in red for positive values, or blue for negative values. Non-significant correlations (p-value $>0.05$) are presented with 0 shown in grey.
@@
The arousal dimension was thought to be more difficult to annotate, as the context of the study, i.e. exercising on a computer and conversing with an AI, is likely to provide little variation for this dimension. This assumption is supported by the fact that the variance of arousal dimension was systematically among the lowest, which was observed for both time-continuous annotations (arousal: .010, novelty: .022, intrinsic pleasantness: .009, goal conduciveness: .017, coping: .020), and summary annotations (arousal: .008, novelty: .026, intrinsic pleasantness: .012, goal conduciveness: .021, coping: .027).
@@ V-C Relation between dimensions and affective labels
We further analysed the agreement between annotators by exploring the relationship between the appraisal dimensions and affective labels. To measure this relationship, we calculated the PCC between each pair of dimension-affective label across all sequences and for each annotator, using the summary values. Then, we evaluated how much the annotators agreed on the correlations between affective labels and dimensions. To better understand this approach, we present as a simplified view in Figure 5 a heatmap of these associations calculated on the average of the evaluations from the six annotators (see section III.B of the supplementary material for a complete view of the heatmap of each annotator). Each cell of the heatmap corresponds to the value of the PCC calculated between a specific dimension (summary) and an affective label.
@@
Some of our results are nonetheless surprising. The anger affective label is correlated with an absence of coping, whereas we would expect the opposite [44], as it generally refers to a situation in which an event is appraised as goal obstructive, but individuals appraise themselves capable of coping with it, e.g., being mugged in the street but wanting to defend themselves. In the context of CCT, where participants sit in front of a computer, different forms of anger may likely occurred, probably involving an inability to cope.
The literature suggests that the affective label of interest would be associated with a balanced mix between coping, i.e., the ability to understand, and novelty [41]. Although our results showed a correlation between interest and coping, an inverse correlation was found between interest and novelty, i.e., higher interest coincided with diminished novelty. This could be explained by the fact that the annotations in our study focused on only one sub-dimension of novelty, i.e., predictability, and did not take into account the other existing sub-dimensions such as complexity and unfamiliarity. As expected, the affective label of surprise is highly correlated with novelty. Even though the surprise affective label is generally considered to be positive or negative depending on the context, it seems to be rather negative in our results as it is correlated with unpleasantness, goal obstructiveness, and no coping. This could be explained by the fact that the occurrence of unexpected situations were more related to negative events such as unexpected negative feedback from the virtual assistant.
Although all the affective labels reported in the Figure 5 seem to be correlated with the appraisal dimensions, with the exception of sad, some correlations actually reflect the result of only one annotator. Affective labels were thus defined as irrelevant in the context AI-assisted CCT if no significant correlation is found with any appraisal dimension for at least two annotators, which was the case for several labels, as reported in Table II; labels consistent with the appraisal dimensions are marked with ✓.
@@ V-D Label core set selection
A core set of the likely most relevant affective labels in the context of AI-assisted CCT was defined based on our analyses. Four main criteria of selection were used: 1) the frequency is significantly higher than the average on all affective labels, 2) the agreement on presence is significantly higher than chance, 3) the agreement on intensity is sufficiently high, and 4) there is no significant correlation with at least one dimension for at least two annotators. Ten affective labels meet these four criteria, five positive and five negative, cf. Table II.
@@ VI Automatic recognition of affective states
All the experiments are based on a single partitioning of the data set that was divided into three partitions: training, validation and test set, consisting of 56 %, 23 % and 21 % of the participants respectively, c.f. Table IV. These partitions were defined to ensure that the statistics – 1) ratio of female vs. male gender, 2) ratio of senior vs MCI participants, 3) ratio of participants based on affective induction, and 4) distribution of education levels – were preserved in the different population groups. All results are reported on the test set.
In the following, we introduce the used representations of the audio, visual, and textual data, along with the predictive models and strategies, and then report on the results.
TABLE IV: Summary statistics of the partitions.
@@ VI-A Representations of audio, textual and video data
We included both expert and deep representations of audio, visual, and textual [^6] In order to make the comparison fair, textual data did not rely on the human transcriptions but on the outputs of an ASR, which is evaluated in details in section IV.A of the supplementary material. data to evaluate their respective contribution in the analysis of affect. Expert based features rely on the statistical description of perceptually salient patterns in the data, and involve relatively light-weight computational resources, whereas deep representations exploit self-supervised learning (SSL) methods on large scale datasets [45, 46, 47].

### v2

@@ IV Data segmentation and annotation
We collected about 80 hours of data from the 52 healthy older partipants, and about 6 hours of data from the 9 MCI participants. We describe in this section the methods used for the segmentation, transcription, and annotation of this data.
@@ IV-A Segmentation, transcription and selection of sequences
The segmentation of the recordings was based on the pragmatic completeness in breath groups, i.e., groups of words that are delimited by pauses taken for breathing, ensuring they formed a single proposition if they conveyed the same semantic information, or separate propositions if they related to different types of semantic information [25]. Non-verbal expressions occurring within or proximate to a proposition were marked by a diacritic, and those occurring at distance were specifically segmented. Temporal markers for proposition start and stop were both positioned at a certain distance to allow a sufficient observation space for the annotation of affect.
The transcription relied mostly on the ESLO project principles [26], respecting spelling and spoken structures; details regarding our conventions are reported in section II.A of the supplementary material. Recordings were transcribed with the ELAN software [27] by five Master students in Language Engineering at the Université Grenoble Alpes with a two-steps process: an initial stage of segmentation and transcription, followed by a subsequent verification stage conducted by another student. The segmented sequences were selected to retain only those containing expressions of affect, and a fraction was randomly removed to reduce the overall duration. Obtained sequences were then all reviewed and some were further removed to reduce redundancy. Statistics of the sequences are given in table I.
@@ IV-B Selection of affective labels and dimensions
The context of CCT aligns closely with the literature on achievement affects [28, 29, 30], which considers the following types of affect:
  1. (i)
    activity affects generated by the task achievement itself;
  2. (ii)
    outcome affects generated by the success or failure results, whether retrospective or prospective;
  3. (iii)
    epistemic affects generated by a form of metacognition on the perception of the achievement;
  4. (iv)
    social affects generated by another individual.
These affects are particularly relevant in the context of AI-assisted healthcare, as they comprehensively cover those that can be experienced during CCT, i.e., those generated by the CCT itself (activity affects), those generated by past or future CCT failure or success (outcome affects), those generated by the perception of the cognitive abilities involved in CCT (epistemic affects), and those generated by AI-patients interactions (social affects). Our categorical representations of affect relied on the following 23 labels: angry, annoyed, anxious, ashamed, confident, contemptuous, curious, desperate, disappointed, embarrassed, excited, frustrated, interested, guilty, happy, hopeful, impatient, proud, relaxed, sad, satisfied, surprised, and upset.
@@ IV-C Annotation tool and guidelines
We took benefit from previous collaborations with the company Viadialog [^2] https://viadialog.com/ engaged in the development of a web-based annotation tool originally tailored for audio data: ANNOT. This tool has been adapted to the needs of our project and is available online [^3] The URL will be provided for camera-ready. The platform allowed annotators – here 6 French students (3 females) from different backgrounds (engineering, business, and biology) and places (Lyon and Toulouse) – to view and annotate the selected audiovisual sequences at home, using their own computer.
The annotation consisted of three steps: a time-continuous annotation, cf. Figure 3, followed by a summary value, for each of the five dimensions, and a summary value of intensity for one or multiple labels. Thus, each sequence was viewed five times, once for each dimension. At the end of the time-continuous annotation, a graph showing the dynamics of the dimension over time was presented to the annotators, who could then validate or perform the annotation again. After the validation, annotators were asked to provide a summary value for the whole sequence. Finally, once all dimensions were annotated, annotators were asked to indicate if one or more labels could describe the affect they had recognised in the sequence and provide an intensity value for them. Annotators could annotate as many labels as desired, or none at all if the sequence lacked any particular affect. Further details on the annotations steps are reported in section II.C of the supplementary material.
@@
Before starting, annotators participated in a training session with the project investigators, including an expert in appraisal theories of emotion. The purpose of this session was to familiarise the annotators with the concepts of appraisal theories, i.e., the meaning of dimensions and affective labels, and to explain in detail the guidelines for the annotation process. Annotators were instructed to put themselves in the participant’s shoes to assess the emotional expression conveyed in the selected sequences, considering factors such as the chosen words, their oral delivery, and the expressive dynamics of the face and body. It was emphasised that there were no right or wrong answers, as the task was inherently subjective. Additionally, they were instructed to perform the task in full-screen mode, in a quiet environment, with sessions lasting no longer than one hour, followed by ten-minute breaks. The annotation process was continuously monitored through an online tracking table, updated daily by the annotators, with additional communication via email or phone as needed. After the training session, each annotator received a document summarising all these points, including definitions of dimensions and labels. Each dimension was defined as follows:
@@ V Corpus Analysis
This section provides a comprehensive statistical analysis of the annotation data. We first assessed the annotation frequency of the affective labels as well as the inter-annotator agreement on their intensity. We then evaluated the cognitive delay involved in the annotation of each dimension, as well as the inter-annotator agreement. Relationships between dimensions and affective labels were also estimated, and a label core set was finally selected, based on compliance with consistency criteria relating to each of the analyses mentioned above.
Figure 4: Frequency (logarithmic scale) of the annotated labels according to the agreement of at least one, two or three annotators among six.
@@ V-A Categorical annotations
The frequency at which the affective labels were selected by at least one, two, or three annotators is reported in Figure 4. Based on the sample of annotation frequency per label for at least one annotator, Cohen’s $d$ was calculated to identify the theoretical value at which the effect size of its deviation from the sample mean can be considered at least "small", i.e., Cohen’s $d$ > 0.2  [33]; this theoretical value was equal to 747 sequences. The annotation frequency per label for at least one annotator is reported in Table II, and shown in bold when above this theoretical value.
@@
where $k$ is the total number of classes, which corresponds here to the presence or absence of a label. $N^{i}$ is the total number of annotated sequences for a given affect label, and $N_{c}^{i}$ is the number of commonly identified classes, whether present or absent. Results are reported in Table II. According to a two-tailored $z$-test, the inter-annotator agreement would be defined as statistically above chance if it would exceed 53%, which is the case for all the labels excepted embarrassed.
TABLE II: Statistics on the annotated affective labels: frequency, inter-annotator agreement on the presence, and intensity, and correlation to appraisal’s dimensions. Most relevant labels in the context of AI-assisted CCT are marked in bold
  Label Frequency Agreement on Cor. to
@@
We further evaluated the inter-rater agreement on the intensity of the affective labels using the Pearson’s Correlation Coefficient (PCC). The inter-rater agreement was defined as sufficiently high if its Cohen’s $d$, i.e., its PCC value divided by the sample standard deviation, is superior to .2 [33], which corresponds to a PCC greater than .065. This concerned most of the affective labels, which are reported in bold style.
TABLE III: Inter-annotator agreement on each sequence, or CCT session, computed with the Pearson’s correlation coefficient for both time-continuous and summary values of appraisal dimensions; mean (standard-deviation)
@@ V-B Dimensional annotations
The cognitive delay in the continuous annotation of each dimension was evaluated, and results showed that goal conduciveness (3.3 seconds) had a significantly longer delay than the other dimensions (less than 2 seconds), suggesting that annotators had more difficulty assessing it. This evaluation is detailed in section III.A of the supplementary material.
@@
We assessed the inter-annotator agreement on the dimensions for both time-continuous and summary annotations, using the PCC, which was computed either on each sequence, or each CCT session, cf. Table III. An high agreement was observed on the dimensions of coping, goal conduciveness, and intrinsic pleasantness, with a more pronounced consistency observed in summary annotations compared to time-continuous annotations, which is expected since time-continuous data imply a greater variation in the possible outcomes. A lower agreement was however found for arousal and novelty, which might be explained by several reasons as detailed below.
@@
The novelty dimension was assumed to be more difficult to annotate, as it probably concerned very specific cases of our study, e.g., unplanned feedback from the virtual assistant on the participant’s performance, occurring in fewer sequences than the other dimensions. This assumption was corroborated by an analysis evidencing that the frequency of high-value annotations was lower for novelty compared to other dimensions; for a high-value threshold set at .85, the proportion of high-value annotations for novelty accounted for only .13% of the time-continuous annotation data, whereas other dimensions were at least twice as frequent: .30% for arousal, .24% for intrinsic pleasantness, .62% for goal conduciveness, and .36% for coping. A similar trend was observed in summary annotations, with the occurrence of high-value annotations for novelty accounting to .26%, whereas all other dimensions were twice as prevalent; arousal: .92%, intrinsic pleasantness: .52%, goal conduciveness: 1.35%, and coping: 2.16%.
Figure 5: Heatmap of the PCC between affective labels and summary values of appraisal dimensions aggregated across all annotators. The sign (+) or (-) after each label denotes whether it is positive or negative, respectively. Dimensions are as follows: AR: arousal, CP: coping, GC: goal conduciveness, IP: intrinsic pleasantness, NV: novelty. The correlations shown represent the average between annotators for whom the correlation was significant (p < 0.05) after Bonferroni correction and of identical direction. Positive or negative direction is shown in blue or red, respectively. A correlation present for a single annotator was considered null.
@@
The arousal dimension was thought to be more difficult to annotate, as the context of the study, i.e. exercising on a computer and conversing with a virtual assistant, is likely to provide little variation for this dimension. This assumption is supported by the fact that the variance of arousal dimension was systematically among the lowest, which was observed for both time-continuous annotations – arousal: .010, novelty: .022, intrinsic pleasantness: .009, goal conduciveness: .017, coping: .020 –, and summary annotations – arousal: .008, novelty: .026, intrinsic pleasantness: .012, goal conduciveness: .021, coping: .027.
@@ V-C Relation between dimensions and affective labels
Further analysis of annotator agreement was conducted by exploring the relationship between summary values of appraisal dimensions and affective labels of each annotator separately. Specifically, the PCC was calculated between each pair of appraisal dimensions and affective labels for all sequences in which the value of the label concerned differed from zero; cf. section III.B of the supplementary material for a comprehensive view of the heatmap of each annotator. A correlation was considered consistent if it was statistically significant and in the same direction for at least two annotators. The average of consistent correlations between annotators are reported in Figure 5. The labels that are consistently correlated with at least one appraisal dimension are marked with "✓" in the Table II.
@@
Some of our results are, however, unexpected. The anger affective label is correlated with an absence of coping, whereas we would expect the opposite [34], as it generally refers to a situation in which an event is appraised as goal obstructive, but individuals appraise themselves capable of coping with it, e.g., being mugged in the street but wanting to defend themselves. In the context of CCT, where participants sit in front of a computer, different forms of anger may likely occur, probably involving an inability to cope. As expected, the affective label of surprise is highly correlated with novelty. Even though the surprise affective label is generally considered to be positive or negative depending on the context, it seems to be rather negative in our results as it is correlated with goal obstructiveness, and no coping. This could be explained by the fact that the occurrence of unexpected situations were more related to negative events such as unexpected negative feedback from the virtual assistant.
@@ V-D Label core set selection
A core set of the likely most relevant affective labels in the context of AI-assisted CCT was defined based on our analyses. Four main criteria of selection were used: (i) the frequency is significantly higher than the average on all affective labels; (ii) the agreement on presence is significantly higher than chance; (iii) the agreement on intensity is sufficiently high; and (iv) there is a significant correlation, in the same direction, with at least one dimension for at least two annotators. Ten affective labels meet these four criteria, five positive and five negative, cf. Table II.
@@ VI Automatic recognition of affective states
All experiments were conducted using a single partitioning of the dataset, divided into three subsets: training, validation and test, comprising 56 %, 23 % and 21 % of the participants respectively; see Table IV. The partitions were carefully designed to minimise bias while maintaining key statistical properties, including: (i) the gender ratio (female vs. male); (ii) the proportion of healthy vs. MCI participants; (iii) the distribution of participants across affective induction conditions; (iv) the spectrum of education levels; and (v) the distribution of label presence, label intensity, and dimensional intensity. Heatmaps depicting these distributions are available in section IV of the supplementary material.
TABLE IV: Summary statistics of the partitions
@@ VI-A Representations of audio, textual and video data
We included both expert and deep representations of audio, visual, and textual [^4] In order to make the comparison fair, textual data did not rely on the human transcriptions but on the outputs of an ASR, which is evaluated in detail in section IV.A of the supplementary material. data to evaluate their respective contribution in the analysis of affect. Expert based features rely on the statistical description of perceptually salient patterns in the data, and involve relatively light-weight computational resources, whereas deep representations exploit self-supervised learning (SSL) methods on large scale datasets [35].